Solve the equation (dy/dx) − y = exy2

1
Expert's answer
2021-09-07T09:56:02-0400

"\frac{dy}{dx} - y = (e^x)y^2" ...(1)

taking "y = \frac{1}{v}"

"\frac{dy}{dx}=-\frac{1}{v^2}.\frac{dv}{dx}"

From (1)

"-\frac{1}{v^2}.\frac{dv}{dx}-\frac{1}{v}=e^x.\frac{1}{v^2}"

"\frac{dv}{dx}+v=-e^x...(2)"

The integrating factor "=e^{\int1dx}=e^x"

Thus, the solution of (2) is

"v.e^x=\int(-e^x).e^xdx+c"

"v.e^x=-\int e^{2x}dx+c"

"v.e^x=-\frac{ e^{2x}}{2}+c"

"\frac{e^x}{y}=-\frac{ e^{2x}}{2}+c"

"\frac{1}{y}=\frac{ -e^{x}+2c.e^{-x}}{2}"

"y=\frac{2}{ -e^{x}+2c.e^{-x}}"